Q12. The following frequency distribution and cumulative frequency distribution represent students grades in a quiz of 10 points. Answer the following Questions: The numter of students who took the quiz. Percent of students who failed the test (less than or equal 60% of the grade) is- Percent of students who passed with better than B grade (better than 80% of the grade) is The majority of students made a grade of FREQUENCY DISTRIBUTION OF STUDENTS GRADES OUT OF 10 2 Cumulative Frequency % 120 100 40 0 1 2 3 78 Grades
Answer #1

a) Number of students = 100

b) Students who fail = 14+14+36 = 64 = 64%

c) Percent with a grade better than 8 = 16/100= 16%

d) majority of sudents have grade between 4 to 6 , ie 36% students.